Still lost, please help!!
Lead pipes and solder are often found in older buildngs. Unfortunately, as lead decays, toxic chemicals can get in the water resting in the pipes. The half-life of lead is 22 yr. What is the decay rate? Thamks

Answer by stanbon(57347)

The half-life of lead is 22 yr. What is the decay rate?
----------
Let "r" be the remainder rate :
----------
Solve: (1/2)Ao = Ao(r)^22)
---
r^22 = (1/2)
r = 0.5^(1/22)
r = 0.97
Decay rate = 1-0.97 = 3%
=============================
Cheers,
Stan H.